### v1

# Identifying non-Hermitian critical points with quantum metric

Abstract
The geometric properties of quantum states is fully encoded by the quantum geometric tensor. The real and imaginary parts of the quantum geometric tensor are the quantum metric and Berry curvature, which characterize the distance and phase difference between two nearby quantum states in Hilbert space, respectively. For conventional Hermitian quantum systems, the quantum metric corresponds to the fidelity susceptibility and has already been used to specify quantum phase transitions from the geometric perspective. In this work, we extend this wisdom to the non-Hermitian systems for revealing non-Hermitian critical points. To be concrete, by employing numerical exact diagonalization and analytical methods, we calculate the quantum metric and corresponding order parameters in various non-Hermitian models, which include two non-Hermitian generalized Aubry-André models and non-Hermitian cluster and mixed-field Ising models. We demonstrate that the quantum metric of eigenstates in these non-Hermitian models exactly identifies the localization transitions, mobility edges, and many-body quantum phase transitions, respectively. We further show that this strategy is robust against the finite-size effect and different boundary conditions.
I Introduction
Quantum phase transitions are a kind of phase transitions that are driven by quantum fluctuations and occur at zero temperature Sachdev (2011). They are extensively investigated between different magnetic phases Smacchia et al. (2011), localization phases Ganeshan et al. (2015); An et al. (2021); Wang et al. (2022); Tang et al. (2021), topological phases Matsumoto et al. (2020); Longhi (2019); Gong et al. (2018); Zhang et al. (2021), and so on. Near the phase transition point, the infinitesimal variation of certain system parameters will dramatically change the properties of the ground (eigen) state, which intuitively leads to a sharp change of the state fidelity. Moreover, an information-geometric approach based on this observation has already been put forward Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010); Ma et al. (2010); Carollo et al. (2020). The concrete procedure is to calculate the fidelity susceptibility, which defines the response of fidelity to the driving parameter of the system Hamiltonian near critical points  Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010).
The fidelity susceptibility is closely related to the quantum metric, which characterizes the distance between two nearby quantum states in Hilbert space Kolodrubetz et al. (2017); Ding et al. (2024, 2022); Törmä (2023). The quantum metric is the real part of a more general geometric concept, called the quantum geometric tensor (QGT) Kolodrubetz et al. (2017); Provost and Vallee (1980), of which the imaginary part is the well-studied Berry curvature Thouless et al. (1982); Simon (1983); Berry (1984) for plenty of celebrated physical effects Guo et al. (2016); Sundaram and Niu (1999); Nagaosa et al. (2010); Aharonov and Bohm (1959); Hasan and Kane (2010); Qi and Zhang (2011); Zhang et al. (2018). The quantum metric is also related to many fascinating physical phenomena, such as the superfluidity in flat bands Julku et al. (2016); He et al. (2021) and topological quantum phases Roy (2014); Lim et al. (2015); Palumbo and Goldman (2018). The quantum metric has recently been experimentally measured in various artificial quantum systems Tan et al. (2019, 2021); Liao et al. (2021a); Chen et al. (2022); Asteria et al. (2019); Yu et al. (2022); Zheng et al. (2022); Lysne et al. (2023); Tan et al. (2018); Yu et al. (2019, 2024); Gianfrate et al. (2020); Yi et al. (2023); Cuerda et al. .
Most of these previous studies focus on the QGT and related quantum phase transitions in Hermitian systems Solnyshkov et al. (2021); Campos Venuti and Zanardi (2007); Zanardi et al. (2007b); Zhang et al. ; Abasto et al. (2008); Albuquerque et al. (2010); Dey et al. (2012). In recent years, the theoretical and experimental explorations of non-Hermitian physics have garnered significant attentions Hatano and Nelson (1996, 1997); Bender and Boettcher (1998); Liu et al. (2019); Kawabata et al. (2019); Chen and Zhai (2018); Zhang et al. (2020a); Hu et al. ; Tu et al. (2023); Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022); Yao and Wang (2018); Jiang et al. (2019); Song et al. (2019); Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019); Li et al. (2023); Liao et al. (2021b). Various novel phenomena or phases unique to non-Hermitian systems have been revealed, such as exceptional points Tzeng et al. (2021); Sticlet et al. (2023); Dóra et al. (2022), non-Hermitian skin effects Yao and Wang (2018); Song et al. (2019); Jiang et al. (2019), and non-Hermitian delocalization Tang et al. (2020); Zhang et al. (2020b); Liu et al. (2020a); Zhai et al. (2020); Liu et al. (2020b); Hamazaki et al. (2019). It is worth mentioning that phase transitions could be also sensitive to the non-Hermiticity, and the associated critical points in non-Hermitian systems have potential for sensitivity enhancement in the quantum sensing regime Li et al. (2023); Liao et al. (2021b); Di Candia et al. (2023); Liang et al. (2020); He et al. (2023).
Here we exploit the information-geometric approach to study quantum phase transitions in both single-particle and many-body non-Hermitian systems. We explore the critical points in various non-Hermitian models through the quantum metric. By using the numerical exact diagonalization, we calculate the quantum metric and localization properties (the fractal dimension and participation ratio) in two non-Hermitian generalized Aubry-André (GAA) models Harper (1955); Aubry and André (1980). Our results show that the quantum metric will be divergent or singular in proximity of localization transition points for all eigenstates. Thus the quantum metric exactly identifies the localization transition and the mobility edges in these non-Hermitian quasiperiodic systems. For non-Hermitian cluster and mixed-field Ising models, we demonstrate that the quantum metric of the ground state can reveal the quantum phase transitions, as compared with the corresponding order parameters obtained from both analytical and numerical approaches. Finally, we show that this information-geometric strategy for identifying non-Hermitian critical points is robust against the finite-size effect and different boundary conditions.
The rest of this paper is organized as follows. A brief review of the quantum metric for non-Hermitian Hamiltonians is given in Sec. II. In Sec. III, we identify the localization transitions and mobility edges in two non-Hermitian GAA models through the quantum metric of eigenstates, respectively. Sec. IV is devoted to revealing quantum phase transitions in non-Hermitian cluster and mixed-field Ising models with the quantum metric of ground states. Brief discussions on the finite-size effect and boundary conditions and a short conclusion are presented in Sec. V.
II Non-Hermitian quantum metric
We begin by considering a non-Hermitian Hamiltonian $H(\bm{\lambda})$ parameterized by $\bm{\lambda}=(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$ in $D$-dimensional parameter space. The eigenstate equation is $H(\bm{\lambda})\ket{\psi_{n}^{R}(\bm{\lambda})}=E_{n}\ket{\psi_{n}^{R}(\bm{\lambda})}$ with $n$ the band index. The eigenvalues $E_{n}$ are in general complex due to the non-Hermiticity of the Hamiltonian, and the eigenstate |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ corresponding to the minimum real eigenvalue is set as the ground state. The eigenstates satisfy the normalization condition $\braket{\psi_{n}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}=1$ but is generally nonorthogonal $\braket{\psi_{m}^{R}(\bm{\lambda})}{\psi_{n}^{R}(\bm{\lambda})}\neq 0$ $(m\neq n)$. We assume the ground state |$\ket{\psi_{1}^{R}(\bm{\lambda})}$ is non-degenerate and focus on the right eigenstates |$\ket{\psi_{n}^{R}(\bm{\lambda})}\doteq\ket{\psi_{n}(\bm{\lambda})}$ for simplicity. The distance between two neighbor eigenstates |$\ket{\psi_{n}(\bm{\lambda})}$ and |$\ket{\psi_{n}(\bm{\lambda}+d\bm{\lambda})}$ in the parameter space is given by Ye et al. ; Provost and Vallee (1980); Kolodrubetz et al. (2017)
$ds^{2}=1-\lvert\braket{\psi_{n}(\bm{\lambda})}{\psi_{n}(\bm{\lambda}+d\bm{\lambda})}\rvert^{2}=\sum_{\mu\nu}g^{(n)}_{\mu\nu}d\bm{\lambda}_{\mu}d\bm{\lambda}_{\nu},$ (1)
where $g^{(n)}_{\mu\nu}$ is the non-Hermitian quantum metric with parameters $\{\mu,\nu\}\in(\lambda_{1},\lambda_{2},\dots,\lambda_{D})$. It is the real part of the non-Hermitian quantum geometric tensor  Ye et al. ; Provost and Vallee (1980); Kolodrubetz et al. (2017)
$\begin{split}Q^{(n)}_{\mu\nu}&=\braket{\partial_{\bm{\lambda}_{\mu}}\psi_{n}}{\partial_{\bm{\lambda}_{\nu}}\psi_{n}}-\braket{\partial_{\bm{\lambda}_{\mu}}\psi_{n}}{\psi_{n}}\braket{\psi_{n}}{\partial_{\bm{\lambda}_{\nu}}\psi_{n}}\\ &=g_{\mu\nu}-\frac{i}{2}F_{\mu\nu},\end{split}$ (2)
where $g_{\mu\nu}^{(n)}=\text{Re}(Q_{\mu\nu})$, and the imaginary part $F_{\mu\nu}=-2\text{Im}(Q_{\mu\nu})$ corresponds to the Berry curvature.
For identifying the quantum phase transition driven by a single parameter $\mu$, we can consider an infinitesimal change of the parameter: $\mu\to\mu+d\mu$, and then obtain
$F_{n}^{2}\doteq\lvert\braket{\psi_{n}(\mu)}{\psi_{n}(\mu+d\mu)}\rvert^{2}\approx 1-\chi^{(n)}_{F}(d\mu)^{2}.$ (3)
Here $F_{n}$ is the so-called fidelity of the $n$-th eigenstate, and $\chi^{(n)}_{F}$ is the fidelity susceptibility Zanardi and Paunković (2006); Zanardi et al. (2007a); Carollo et al. (2020). Notably, $\chi^{(n)}_{F}$ denotes the response of $F_{n}$ due to the small change of $\mu$. Combining Eq. (1) and Eq. (3), we can proof the equivalence of the quantum metric and the fidelity susceptibility in generic non-Hermitian Hamiltonians:
$\begin{split}g^{(n)}_{\mu\mu}&=\braket{\partial_{\mu}\psi_{n}}{\partial_{\mu}\psi_{n}}-\braket{\partial_{\mu}\psi_{n}}{\psi_{n}}\braket{\psi_{n}}{\partial_{\mu}\psi_{n}}\\ &=\lim_{d\mu\to 0}\frac{-2\ln F_{n}}{d\mu^{2}}=\chi^{(n)}_{F}.\end{split}$ (4)
It has been shown that the fidelity susceptibility of the ground state can detect quantum critical points in Hermitian systems  Zanardi and Paunković (2006); Zhu (2006, 2008); Zanardi et al. (2007a); You et al. (2007); Gu (2010). For $n$-th eigenstate, one can also expect that the corresponding quantum metric $g^{(n)}_{\mu\mu}$ will reach the maximum value or diverge at its localization transition point. In the following, we extend this wisdom to generic non-Hermitian models. To be concrete, we demonstrate that the localization transition of all eigenstates with related mobility edges and the quantum phase transitions in non-Hermitian disordered and many-body systems can be revealed by the quantum metric, respectively.
III Non-hermitian localization transitions and mobility edges
III.1 Detecting localization transitions
Figure 1: (Color online) The non-Hermitian GAA model with nonreciprocal hopping given by $H_{\text{GAA}}^{(1)}$ with $L=610$ under the PBC. (a) $\eta$ and (b) $\xi_{V_{2}}$ for the ground state in the parameter space ($V_{1}$,$V_{2}$). The localization phase diagram in (a) consists three regions $\mathrm{I}$, $\mathrm{II}$, $\mathrm{III}$ for the extended, critical, and localized phases, respectively. (c,d) $\xi_{V_{1}}$ and ${\eta}$ as a function of $V_{1}$ with fixed $V_{2}=0.5$ and $V_{2}=1.5$, respectively. (e) $\xi_{V_{2}}$ and ${\eta}$ as a function of $V_{2}$ with fixed $V_{1}=3$. (f) $\xi_{V_{1}}$ as functions of $V_{1}$ and $g$ with fixed $V_{2}=0.5$. The colored phase regions with boundaries in (c-e) are determined by Eq. (8). The nonreciprocal strength is $g=0.5$ in (a-e). Other parameters are $h=0$ and $t=1$.
Figure 2: (Color online) The non-Hermitian GAA model with complex potential phase (and nonreciprocal hopping) given by $H_{\text{GAA}}^{(1)}$ with $L=610$ under the PBC. (a) $\eta$ and (b) $\xi_{V_{1}}$ in the parameter space ($V_{1}$,$V_{2}$) for $g=0$. The three regions $\mathrm{I}$, $\mathrm{II}$, and $\mathrm{III}$ denotes the extended, critical, and localized phases, respectively. (c) $\xi_{V_{1}}$ and ${\eta}$ as a function of $V_{1}$ with fixed $V_{2}=0.5$ and $g=0.5$. (d) $\xi_{V_{2}}$ and ${\eta}$ as a function of $V_{2}$ with fixed $V_{1}=1.8$ and $g=0.5$. The colored phase regions with boundaries in (c,d) are determined by Eq. (8). Other parameters are $h=0.5$ and $t=1$.
We first consider a non-Hermitian version of the GAA model with quasi-periodic modulations on both the on-site potential and off-diagonal hopping amplitude Chang et al. (1997); Liu et al. (2015); Wang et al. (2021). Generally, the non-Hermiticities come from the nonreciprocal hopping and complex potential phase, and the model Hamiltonian can be written as Tang et al. (2021); Harper (1955); Aubry and André (1980)
$H_{\text{GAA}}^{(1)}=\sum_{j}[t_{j}(e^{-g}c^{\dagger}_{j+1}c_{j}+e^{g}c^{\dagger}_{j}c_{j+1})+\Delta_{j}c^{\dagger}_{j}c_{j}],$ (5)
where $c_{j}$ ($c_{j}^{\dagger}$) is the annihilation (creation) operator at the $j$-th site. Here $g$ is the dimensionless nonreciprocal strength of the hopping amplitude $t_{j}$, and $\Delta_{j}$ is the on-site potential term. They are given by
$\begin{split}t_{j}&=t+V_{2}\cos[2\pi\beta(j+1/2)],\\ \Delta_{j}&=V_{1}\cos(2\pi\beta j+ih),\\ \end{split}$ (6)
with $V_{1}$ and $V_{2}$ being the modulation amplitudes of the potential and hopping terms, respectively, $h$ being the complex phase, and $\beta$ being an irrational number to ensure the incommensurate modulation. Hereafter, we choose $\beta=(\sqrt{5}-1)/2$ as the golden ratio and the lattice site number $L$ from the Fibonacci sequence, and set $t=1$ as the unit of energy.
The GAA model has rich localization phase diagrams in the Hermitian Chang et al. (1997); Liu et al. (2015); Wang et al. (2021) and non-Hermitian cases Tang et al. (2021). By changing the modulation amplitudes $V_{1}$ and $V_{2}$ (or $g$), the localization transition of eigenstates between extended, localized, and critical phases can occurs. A typical quantity to characterize localization properties of the $n$-th eigenstate in this model is the fractal dimension:
$\eta_{n}=-\frac{\ln(\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{2})}{\ln L},$ (7)
where |$\ket{j}$ is the computational basis and |$\ket{\psi_{n}}$ is the $n$-th right eigenstates. In this model, there is no mobility edge and all eigenstates share the same localization properties Chang et al. (1997); Liu et al. (2015); Wang et al. (2021); Tang et al. (2021). Thus, we focus on the ground state |$\ket{\psi_{1}}$ with the corresponding fractal dimension $\eta_{1}\equiv\eta$ for simplicity. When the system size $L\to\infty$, one has $\eta\to 1$ for the extended phase, $\eta\to 0$ for the localized phase, while $\eta$ falls between 00 and $1$ ($\eta\approx 0.5$) in the critical phase, respectively. Figure 1(a) show the numerical results of $\eta$ in the parameter space ($V_{1}$,$V_{2}$) for nonreciprocal case with $g=0.5$ and $h=0$ under the periodic boundary condition (PBC), by using the exact diagonalization method. Three regimes (denoted by the I, II and III) for the extended, critical, and localized phases are exhibited. Notably, the critical values $V_{1c}$ for localization transition points with respect to other parameters are given by Tang et al. (2021)
$V_{1c}=e^{-|h|}\left(2K\cosh|g|+2\sqrt{K^{2}-V_{2}^{2}}\sinh|g|\right),$ (8)
with $K=\max(t,V_{2})$.
To demonstrate the ability of the non-Hermitian quantum metric to detect localization transition points in non-Hermitian cases, we numerically calculate the diagonal elements of the quantum metric $g_{\mu\mu}$ for the parameter $\mu=\{V_{1},V_{2}\}$. To resole the three different localization regions in this model, we use the logarithm quantity of the quantum metric
$\xi_{\mu}=\log_{10}g_{\mu\mu},$ (9)
with respect to the model parameters $(V_{1},V_{2},g,h)$. We first consider the generalized AA model with only nonreciprocal hopping for fixed $h=0$ and $g=0.5$. The numerical result of $\xi_{V_{2}}$ in Fig. 1(b) shows the boundaries between different localization phase regions, which is consistent with those in Fig. 1(a). To reveal the localization transition points more clearly, we plot $\xi_{V_{1}}$ and $\eta$ as a function of $V_{1}$ for $V_{2}=0.5$ and $V_{2}=1.5$ in Fig. 1(c) and Fig. 1(d), respectively. The results show that $\xi_{V_{1}}$ changes relatively smoothly in each phase, while increases significantly and reaches its maximum value at the localization transition points. In Fig. 1(e), we plot $\xi_{V_{2}}$ and $\eta$ as a function of $V_{2}$ for $V_{1}=3$, which across three localization phases. In this case, we can see that the boundaries between different phases can also be revealed, although the rapid oscillation of $\xi_{V_{2}}$ in the critical phase region is exhibited due to the accidental degeneracies of wave functions with respect to $V_{2}$ in the critical phase. The quantum metric can even be used to identify the localization transition driven by the non-Hermitian strength $g$, as shown in Fig. 1(f).
We also consider the GAA model with the complex potential phase for $h=0.5$ and $g=0$, with numerical results of $\eta$ and $\xi_{V_{1}}$ shown in Fig. 2(a) and Fig. 2(b), respectively. It is evident that the boundaries between the three localization phases distinguished by $\eta$ corresponds to the peak of $\xi_{V_{1}}$. For the case of coexistence of the nonreciprocal hopping and complex potential phase with $g=h=0.5$, one can still find the peaks of the quantum metric ($\xi_{V_{1}}$ and $\xi_{V_{2}}$) at the localization transition points in Fig. 2(c) and Fig. 2(d). Moreover, we plot three phase regions with different colors in Figs. 1(c-e) and Figs. 2(c,d) according to Eq. (8). One can see that localization transition points in this non-Hermitian GAA model revealed by the quantum metric are perfectly consistent with those obtained from the fractal dimension and the analytical result Tang et al. (2021).
III.2 Revealing mobility edges
Figure 3: (Color online) The GAA model given by $H_{\text{GAA}}^{(2)}$ in the Hermitian limit with $L=610$ under the PBC. The eigenenergies $E_{n}$ as functions of (a) $\Delta$ and (c) $\alpha$, respectively. The colors correspond to the values of $\mathrm{PR}_{n}$ as the inset color bar. The red circles denote the positions of the peaks of the quantum metric for eigenstates at different $\Delta$ or $\alpha$. The black solid line denotes the exact mobility edge given by Eq (12). Note that $\mathrm{PR}_{n}=0.05$ is set as the critical value for the comparison with the exact mobility edge. (b) $\mathrm{PR}_{n}$ and $g_{\Delta\Delta}^{(n)}$ as functions of $\Delta$ for the ground state ($n=1$) and the highest excited state ($n=L$) with fixed $\alpha=-0.5$. (d) $\mathrm{PR}_{n}$ and $g_{\alpha\alpha}^{(n)}$ as functions of $\alpha$ for the ground state and the highest excited state with fixed $\Delta=1.8$. Other parameters are $g=0$ and $t=1$.
Figure 4: (Color online) The non-Hermitian GAA model given by $H_{\text{GAA}}^{(2)}$ with $L=610$ under the PBC. The eigenenergies $E_{n}$ as functions of (a) $\Delta$ and (c) $\alpha$, respectively. The red circles denote the positions of the peaks of the quantum metric for eigenstates at different $\Delta$ or $\alpha$. (b) $\mathrm{PR}_{1}$ and $g_{\Delta\Delta}^{(1)}$ as a function of $\Delta$ for the ground state with fixed $\alpha=-0.5$. (d) $\mathrm{PR}_{1}$ and $g_{\alpha\alpha}^{(1)}$ as a function of $\alpha$ for the ground state with fixed $\Delta=1.8$. Other parameters are $g=0.5$ and $t=1$.
We proceed to consider localization transitions of all eigenstates to reveal the mobility edge in non-Hermitian systems with the quantum metric. To do this, we study another non-Hermitian version of the GAA model Ganeshan et al. (2015); An et al. (2021); Wang et al. (2022):
$H_{\text{GAA}}^{(2)}=\sum_{j}[t(e^{-g}c^{\dagger}_{j+1}c_{j}+e^{g}c^{\dagger}_{j}c_{j+1})+\epsilon_{j}c^{\dagger}_{j}c_{j}],$ (10)
with a quasi-periodic onsite potential
$\epsilon_{j}=\frac{\Delta\cos(2\pi j\beta)}{1-\alpha\cos(2\pi j\beta)}.$ (11)
Here $g$ represents the nonreciprocal strength of the hopping amplitude, $\Delta$ is the amplitude of irrational modulation with $\beta=(\sqrt{5}-1)/2$, and $\alpha\in(-1,1)$ denotes the lattice parameter. In the Hermitian case ($g=0$), when $\alpha=0$, it reduces to the standard AA model without mobility edges. When $\alpha\neq 0$, this GAA model in the Hermitian limit has an exact mobility edge at energy $E=E_{c}$ that follows the relationship Ganeshan et al. (2015)
$E_{c}=\frac{2t-\Delta}{\alpha}.$ (12)
The presence of the mobility edge can be understood by the energy-independent localization transition and self-duality condition driven by the cosine dispersion Ganeshan et al. (2015). This GAA model has been realized with ultracold atoms in tunable synthetic lattices and the mobility edge has been observed An et al. (2021); Wang et al. (2022). However, the mobility edge in the non-Hermitian GAA model is yet to be explored.
To characterize the localization properties of all eigenstates and reveal the mobility edge in the Hermitian and non-Hermitian cases, we can numerically compute the participation ratio of the $n$-th eigenstate $|\psi_{n}\rangle$:
$\mathrm{PR}_{n}=\frac{1}{L}\frac{1}{\sum_{j}^{L}|\braket{j}{\psi_{n}}|^{2}}.$ (13)
The extended and localized eigenstates take $\mathrm{PR}_{n}\to 1$ and $\mathrm{PR}_{n}\to 0$, respectively. To detect the mobility edge with the quantum metric, we numerically calculate the corresponding diagonal element of the quantum metric $g^{(n)}_{\mu\mu}$ with $\mu=\{\Delta,\alpha\}$ for the $n$-th eigenstate $|\psi_{n}\rangle$.
We first show the numerical results in the Hermitian limit with $g=0$ in Fig. 3. As shown in Fig. 3(a), the mobility edge exhibits and separates the eigenstates from $\mathrm{PR}_{n}$: the upper ones as extended states while the lower ones are localized states. The peak positions of $g^{(n)}_{\Delta\Delta}$ and the exact mobility edge given by Eq. (12) are plotted, which are consistent with each other. To be more clearly, we show $g^{(n)}_{\Delta\Delta}$ and $\mathrm{PR}_{n}$ as a function of $\Delta$ for the ground state ($n=1$) and the highest excited state ($n=L$) in Fig. 3(b), where $g^{(n)}_{\Delta\Delta}$ exhibits a distinct peak at the critical point with $\mathrm{PR}_{n}\approx 0$. The mobility edge separating extended and localized eigenstates with respect to varying the parameter $\alpha$ can be also revealed by the quantum metric $g^{(n)}_{\alpha\alpha}$, as shown in Figs. 3(c,d).
For the non-Hermitian case, the nonreciprocal strength $g$ modify the mobility edge from the exact form, with two examples of $g=0.5$ shown in Fig. 4. In these cases, the localization transitions of right eigenstates with respect to the parameters $\Delta$ ($\alpha$) are revealed by $g^{(n)}_{\Delta\Delta}$ ($g^{(n)}_{\alpha\alpha}$) in Fig. 4(a) [4(c)]. Figure  4(b) [4(d)] shows $\mathrm{PR}_{1}$ and $g^{(1)}_{\Delta\Delta}$ ($g^{(1)}_{\alpha\alpha}$) as a function of $\Delta$ ($\alpha$) for the ground state. One can find the peak of $g^{(1)}_{\Delta\Delta}$ ($g^{(1)}_{\alpha\alpha}$) is exhibited at the localization transition point. Comparing with the Hermitian case, the mobility edge with respect to $\Delta$ moves to the larger values of $\Delta$ in Fig 4(a), while it moves towards larger $|\alpha|$ in Figs 4(c). This can be understood by the delocalization enhanced by the nonreciprocal hopping Hatano and Nelson (1996, 1997); Gong et al. (2018). Thus, our results demonstrate the quantum metric can be used to reveal the mobility edge in both the Hermitian and non-Hermitian systems.
IV Quantum phase transitions in non-hermitian many-body systems
IV.1 Non-Hermitian cluster Ising model
We proceed to reveal the quantum phase transition of many-body ground states in non-Hermitian interacting spin systems with the quantum metric. We first consider the non-Hermitian cluster Ising model with exactly solvable ground states Guo et al. (2022), which takes the following Hamiltonian
$H_{\text{NH-Ising}}^{(1)}=-J\sum_{l=1}^{N}\sigma_{l-1}^{x}\sigma_{l}^{z}\sigma_{l+1}^{x}+\lambda\sum_{l=1}^{N}\sigma_{l}^{y}\sigma_{l+1}^{y}+\frac{i\Gamma}{2}\sum_{l=1}^{N}\sigma_{l}^{u},$ (14)
where $\sigma_{l}^{x,y,z}$ denote the Pauli matrices of the $l$-th spin, and the gain and loss is given by $\sigma_{l}^{u}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$. There are three parameters in this model: $J$ represents the cluster exchange strength, $\lambda$ denotes the Ising exchange strength, and $\Gamma$ signifies the strength of gain or loss. In the rest of this section, we set $J=1$ as the energy unit.
Using the standard Jordan-Wigner transformation and the Fourier transformation under the PBC, we can obtain the Hamiltonian in momentum space, which is given by
$H_{k}=2\sum_{k>0}[iy_{k}(c_{k}^{\dagger}c_{-k}^{\dagger}+c_{k}c_{-k})+z_{k}(c_{k}^{\dagger}c_{k}+c_{-k}^{\dagger}c_{-k}-1)],$ (15)
here $y_{k}=\sin(2k)+\lambda\sin(k)$ and $z_{k}=\cos(2k)-\lambda\cos(k)-\frac{i\Gamma}{4}$. By diagonalizing $H_{k}$ through the Bogoliubov transformation, one can obtain the ground state of this model
|$\ket{G}=\prod_{k>0}[\cos(\frac{\theta_{k}}{2})+i\sin(\frac{\theta_{k}}{2})c_{k}^{\dagger}c_{-k}^{\dagger}]\ket{\mathrm{Vac}},$ (16)
where |$\ket{\mathrm{Vac}}$ is the vacuum state of the free fermion and $\theta_{k}=\arctan(-y_{k}/z_{k})$. The properties of the ground state can be characterized by the staggered magnetization $m_{y}$ and the string order parameter $O^{x}$. The staggered magnetization $m_{y}$ is determined by the spin-correlation function $R_{r}^{y}$ with the following relationship
$m_{y}^{2}=\lim_{r\to\infty}(-1)^{r}R_{r}^{y}=\lim_{r\to\infty}(-1)^{r}\braket{G}{\sigma_{l+r}^{y}\sigma_{l}^{y}}{G},$ (17)
where $r$ is the relative distance between spins and $l$ ($l\in[1,N]$) is an irrelevant index under the PBC. Using the Wick theorem Wick (1950), $R_{r}^{y}$ can be transformed to a Toeplitz determinant
$R^{y}_{r}=\begin{vmatrix}D_{1}&D_{0}&\dots&D_{-r+2}\\ D_{2}&D_{1}&\dots&D_{-r+3}\\ \vdots&\vdots&\ddots&\vdots\\ D_{r}&D_{r-1}&\dots&D_{1}\end{vmatrix},$ (18)
with the elements
[…]
The second order parameter is the string order parameter $O^{x}$, which is given by
$O^{x}=\lim_{r\to\infty}(-1)^{r}\langle G|\sigma_{1}^{x}\sigma_{2}^{y}(\prod_{l=3}^{r}\sigma^{z}_{l})\sigma_{r+1}^{y}\sigma_{r+2}^{x}|G\rangle.$ (20)
Using the same method in Ref. Wick (1950), one can rewrite $O^{x}$ as
$O^{x}=\lim_{r\to\infty}\begin{vmatrix}G_{-2}&G_{-3}&\dots&G_{-r-1}\\ G_{-1}&G_{-2}&\dots&G_{-r}\\ G_{0}&G_{-1}&\dots&G_{-r+1}\\ \vdots&\vdots&\ddots&\vdots\\ G_{r-2}&G_{r-5}&\dots&G_{-3}\\ G_{r-3}&G_{r-4}&\dots&G_{-2}\end{vmatrix},$ (21)
[…]
Figure 5: (Color online) The non-Hermitian cluster Ising model $H_{\text{NH-Ising}}^{(1)}$ under the PBC. (a) The phase diagram with boundaries (blue solid lines) determined by the order parameters $O^{x}$ and $m_{y}$. Four regions $\mathrm{I}$, $\mathrm{II}$, $\mathrm{III}$, $\mathrm{IV}$ denote the cluster, gapless, paramagnetic, and antiferromagnetic phases, respectively. The locations of the peaks of the quantum metric are plotted as the red circles and consistent with the phase boundaries. (b) $g_{\Gamma\Gamma}$ as a function of $\Gamma$ with fixed $\lambda=0.5$ and $\lambda=0.9$. (c) In the Hermitian limit with $\Gamma=0$ and (d) in the non-Hermitian case with $\Gamma=2$, $g_{\lambda\lambda}$, $O^{x}$, and $m_{y}$ as a function of $\lambda$. Other parameter is $J=1$.
By numerically calculating the order parameters $O^{x}$ and $m_{y}$ for $r=1000$ that is large enough to approach the thermodynamic limit, one can obtain four phases in the phase diagram in the $\lambda$-$\Gamma$ plane Guo et al. (2022), as shown in Fig. 5(a). The first one is the gapped cluster phase with long-range correlation ($O^{x}\neq 0$) and vanishing staggered magnetization ($m_{y}=0$). The second one is the gapless phase with the power-law decay of correlation ($O^{x}\neq 0$ for finite systems) and vanishing staggered magnetization ($m_{y}=0$). The third and fourth ones are the conventional paramagnetic ($m_{y}=0$) and antiferromagnetic ($m_{y}\neq 0$) without long-range correlation, respectively. The phase boundaries determined by the order parameters are plotted as the blue solid lines in Fig. 5(a). To show that the quantum phase transition points in this non-Hermitian cluster Ising model can be revealed by the quantum metric, we plot the locations of the peaks of $g_{\mu\mu}$ of the ground state |$\ket{G}$ for the parameter $\mu=\{\lambda,\Gamma\}$. As shown in Fig. 5(a), the locations of the peaks correspond to the critical points of the four different phases. To illustrate more clearly, we plot $g_{\Gamma\Gamma}$ as a function of $\Gamma$ in Fig. 5(b). One can still find the peaks at $\Gamma\approx 2.23$ with fixed $\lambda=0.5$, and at $\Gamma\approx 0.48$ and $\Gamma\approx 2.95$ with fixed $\lambda=0.9$ in $g_{\Gamma\Gamma}$, indicating one and two quantum critical points, respectively. We further plot $g_{\lambda\lambda}$ as a function of $\lambda$ in the Hermitian ($\Gamma=0$) and non-Hermitian case ($\Gamma=2$) in Fig. 5(c) and Fig. 5(d), respectively. The results show that the values of $g_{\lambda\lambda}$ exhibit peaks at the quantum phase transition points where the behaviours of the order parameters $O^{x}$ and $m_{y}$ change.
Figure 6: (Color online) The non-Hermitian mixed-field Ising model $H_{\text{NH-Ising}}^{(2)}$ with the system size $N=10$. (a) $|M_{z}|$ and (b) $\xi_{h_{z}}$ of the ground state as functions of $h_{z}$ and $h_{x}$ under periodic boundary condition. The phase diagram based on $|M_{z}|$ consists two regions $\mathrm{I}$ and $\mathrm{II}$ for the paramagnetic and ferromagnetic phases with real and complex energies, respectively. $|M_{z}|$ and $\xi_{h_{z}}$ as a function of $h_{z}$ (c) under the PBC; and (d) the OBC, respectively. Other parameter is $J=1$.
IV.2 Non-Hermitian mixed-field Ising model
We further study the non-Hermitian mixed-field Ising model, which is non-integrable and described by the following Hamiltonian
$H_{\text{NH-Ising}}^{(2)}=-J\sum_{l=1}^{N}\sigma_{l}^{z}\sigma_{l+1}^{z}+h_{x}\sum_{l=1}^{N}\sigma_{l}^{x}+ih_{z}\sum_{l=1}^{N}\sigma_{l}^{z},$ (22)
where $h_{x}$ and $h_{z}$ denote the strengths of the real and imaginary fields that are along the transverse and longitudinal directions, respectively. The order parameter for the ground state $|\psi_{1}\rangle$ in this model is the magnetization $M_{z}=\frac{1}{N}\bra{\psi_{1}}\sum_{l=1}^{N}\sigma^{z}_{l}\ket{\psi_{1}}$, where |$\ket{\psi_{1}}$ can be numerically obtained by the exact diagonalization method for small system size $N$. The numerical results for $N=10$ under the PBC and the open boundary condition (OBC) are shown in Figs. 6(a-c) and Fig. 6(d), respectively. The phase diagram that contains the paramagnetic phase with $|M_{z}|\approx 0$ (and real energy) and the ferromagnetic phase with finite value of $|M_{z}|$ (and complex energy) in the $h_{z}$-$h_{x}$ parameter plane is shown in Fig. 6(a). Moreover, the phase transition between two phases is accompanied by a real-complex transition of the ground state energy. To detect the phase transition with the quantum metric, we numerically compute the logarithm quantify of quantum metric $\xi_{h_{z}}=\log_{10}g_{h_{z}h_{z}}$ as a function of $h_{z}$ and $h_{x}$ in Fig 6(b). One can see the boundary between the paramagnetic and ferromagnetic phases are well revealed by $\xi_{h_{z}}$. As shown in Fig. 6(c) and Fig. 6(d) for different boundary conditions with fixed $h_{x}=3$, the critical point $h_{z}\approx 0.9$ and $h_{z}\approx 0.96$ can both be clearly revealed by the quantum metric.
V discussion and conclusion
Figure 7: (Color online) (a) $\xi_{V_{1}}$ as a function of $V_{1}$ for $H_{\text{GAA}}^{(1)}$ with different system sizes $L$ and fixed $g=h=0$ under periodic boundary condition. $\xi_{V_{1}}$ as a function of $V_{1}$ for $\tilde{H}$ [see Eq. (23)] with fixed (b) $g=0$; and (c) $g=0.5$, for different boundary condition coefficients $\zeta$. (d) The probability amplitudes of the ground-state wave functions as a function of $V_{1}$ under the PBC and the OBC, respectively. The system size in (b-d) is $L=610$. Other parameters are $h=0$, $V_{2}=0.5$, and $t=1$.
Before concluding, we discuss the finite-size effect and boundary conditions on detecting the non-Hermitian critical points with the quantum metric. Without loss of generality, we focus on the non-Hermitian GAA model $H_{\text{GAA}}^{(1)}$ in Eq. (5). The finite-size scaling of the quantum metric of the ground state $\xi_{V_{1}}=\log_{10}g_{V_{1}V1}$ under the PBC for the lattice sizes $L=34,144,610,2584$ (chosen from the Fibonacci sequences) is shown in Fig 7(a). One can find that the critical point can be revealed by the peak of the quantum metric even for a small lattice of $L=34$, and the peak becomes higher as the lattice size is increased. To consider the boundary effect on the quantum metric, we add an boundary coupling term to the Hamiltonian $H_{\text{GAA}}^{(1)}$ as
$\begin{split}\tilde{H}=H_{\text{GAA}}^{(1)}+\xi(e^{-g}c^{\dagger}_{1}c_{L}+e^{g}c^{\dagger}_{L}c_{1}),\end{split}$ (23)
where $\zeta$ is the coefficient for tuning different boundary conditions. For $\zeta=0$, it denotes the open boundary condition, while $\zeta=1$ ($\zeta\neq 0,1$) denotes the standard (modified) periodic boundary condition. As shown in Fig. 7(b), the results of $\xi_{V_{1}}$ as a function of $V_{1}$ under different boundary conditions ($\zeta=0,0.5,1$) coincide in the Hermitian limit. In the non-Hermitian case, the results of $\xi_{V_{1}}$ are different for periodic ($\zeta=0.5,1$) and open ($\zeta=0$) boundary conditions, while the transition point can revealed in both conditions, as shown in Fig 7(c). The difference comes from the localized bulk eigenstates induced by the non-Hermitian skin effect under the open boundary condition Yao and Wang (2018); Song et al. (2019), as shown in Fig 7(d). The non-Hermitian skin effect destroys the extended phase under the original periodic boundary condition and leads to the increase of the quantum metric with accidental degeneracies. Thus, our results are robust against to the finite-size effect and different boundary conditions when using the quantum metric to reveal critical points.
In summary, we have investigated the quantum phase transitions in various non-Hermitian systems by using the information-geometric approach. It has been shown that the peaks of the quantum metric of the eigenstates exactly identify the localization transition points and mobility edges in the non-Hermitian GAA models. For non-Hermitian cluster and mixed-field Ising models, we have demonstrated that the phase boundaries in the non-Hermiticity parameter space determined by quantum metric of ground state perfectly coincide with those from the corresponding order parameters. These results indicate that the peak of the quantum metric serving as good signatures for detecting the non-Hermitian critical points in both single-particle and many-body systems. This strategy is robust against the finite-size effect and different boundary conditions. Notably, our present work only focuses on the right eigenstates of non-Hermitian Hamiltonians, and the quantum metric effect based on the biorthogonal eigenstates and the non-unitary dynamics are deserving to be further explored.
Acknowledgements.
We thank Zhi Li for helpful discussions. This work was supported by the National Natural Science Foundation of China (Grant No. 12174126), the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024B1515020018), and the Science and Technology Program of Guangzhou (Grant No. 2024A04J3004).